OpenStudy (anonymous):

Medal and fan!!!! 16!=

mathslover (mathslover):

\(n! = n ( n-1) (n-2) ... (3)(2)(1) \) So, 16! = 16 ( 16-1) (16-2) (16-3) ... (3)(2)(1)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=16!

OpenStudy (anonymous):

would it go 16x15x14x13x12x11x10x9x8dx7x6x5x4x3x2x1

mathslover (mathslover):

Yes Punkchick

OpenStudy (anonymous):

really big number use a calculator yes, that is it

mathslover (mathslover):

Yeah, its really bignumber...